OpenStudy (anonymous):

A car company tested a sports car on a road with different inclines. The test driver tested the car by driving a distance of x miles on a flat road, (x2 + 3) miles downhill, and (x - 7) miles uphill. Which simplified expression is equivalent to the total distance, in miles, for which the car was tested?

OpenStudy (anonymous):

\[x+\left(x^2+3\right) +(x-7) =x^2+2 x-4 \]

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

Tnk u so much :) i appreciate it ! x3

OpenStudy (anonymous):

NP